Question7 1 pts You want to buy a new surfboard 4 years from now. You have $500 in the bank today. You can earrn 11% on your savings. If the surfboard will cost $5,000, how much will you have to save every year to have enough to buy it? Round your answer to 2 decimals, for example 100.12. Question 8 1 pts Today you purchased XYZ stock for $92 per share. You will keep the stock for 10 years and then sell it. You hope to make an average annual return of at least 8% on this investment. What does the share price have to be in the future for you to make the desired rate of return? Round your answer to 2 decimals, for example 100.12. Question 9 1 pts You are thinking of buying a new car for $20,000. You will borrow the money to buy the car, and payback the loan in annual payments over 5 years. If the interest rate is 9% per year, what is your payment amount? Round your answer to 2 decimals, for example 100.12.
Answer #1

7.

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

Hence future value of $500=$500*(1.11)^4

Also:

Future value of annuity=Annuity[(1+rate)^time period-1]/rate

=Annuity[(1.11)^4-1]/0.11

Hence

5000=$500*(1.11)^4+Annuity[(1.11)^4-1]/0.11

5000=759.035205+Annuity*4.709731

Hence annuity=(5000-759.035205)/4.709731

=$900.47(Approx).

8.

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.
Hence

A=$92*(1.08)^10

=$92*2.158924997

=$198.62(Approx).

9.

Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

20000=Annuity[1-(1.09)^-5]/0.09

20000=Annuity*3.889651263

Annuity=20000/3.889651263

=$5141.85(Approx).